# Asymptotic stability equals exponential stability—while you twist your eyes

### v2

@@ 1   Introduction
In the context of what we call today Conley index theory (see Section A), Conley posed the following “converse question” in the 1970s: “To what extent does the homotopy index [Conley index] itself determine the equivalence class of isolated invariant sets which are related by continuation?” [ref:conley1978isolated, p. 83]. Recently, Kvalheim proved that uniquely integrable $C^{0}$ vector fields, on a $C^{\infty}$ manifold $M$, rendering a compact set $A\subseteq M$ asymptotically stable, are homotopic on an open neighbourhood $U\supseteq A$ such that throughout the homotopy the vector fields do not vanish on $U\setminus A$ [ref:kvalheim2022obstructions, Thm. 1]. Connecting this result to Conley, a follow-up question (revitalized) by Kvalheim—which is the central question of this note—is the following.
@@
This question, in one form or another, inspired several works, for instance, [ref:reineck1992continuation, ref:mrozek2000conley, ref:JongeneelSchwan2024TAC]. Here, we will elaborate on commentary by the author in [ref:JongeneelSchwan2024TAC]. Specifically, in the seminal paper “Asymptotic stability equals exponential stability, and ISS equals finite energy gain—if you twist your eyes” from 1999, Grüne, Sontag and Wirth showed that asymptotic stability “equals” exponential stability in the sense that if an equilibrium point is asymptotically stable under some vector field on $\mathbb{R}^{n}$, then, there is a suitable change of coordinates rendering this point exponentially stable [ref:grune1999asymptotic]. Such a change of coordinates is understood to be instantaneous. However, by leveraging their work, we show that asymptotic stability can be continuously “transformed” into exponential stability, while preserving asymptotic stability throughout the transformation, see Theorem 3.6. Differently put, asymptotic stability equals exponential stability—not only if you twist your eyes, but while you twist your eyes. The key to this is to observe that we can select the transformation from [ref:grune1999asymptotic] to be an orientation-preserving homeomorphism on $\mathbb{R}^{n}$, not just any homeomorphism cf[ref:JongeneelSchwan2024TAC, Sec. III], see the proof of Proposition 3.2.
@@ 1.1   Related work
It can be argued that questions of the form above emerged from studies aimed at classifying manifolds, maps and so forth. A successful, yet coarse, resolution has been found in the study of these objects up to homotopy, e.g., motivated early on by the fundamental group being homotopy invariant, Hopf’s degree theorem, CW complexes, intractability of topological equivalence and more at the intersection of topology and dynamical systems.
@@
Besides possibly stronger necessary conditions for continuous stabilization, the study of Conley’s converse question directly relates to understanding topological properties of spaces of stable dynamical systems (as was motivation of prior work [ref:JongeneelSchwan2024TAC]). This is of practical importance as several frameworks in optimal control and reinforcement learning aim to optimize over precisely such a space, e.g., moving from a pre-stabilized system to an optimally controlled system, we point to [ref:Fazel_18, ref:furieri2024learning] and references therein.
@@
$[0,1]\ni s\mapsto A(s):=s\cdot\begin{pmatrix}-1&10\\ 0&-1\end{pmatrix}+(1-s)\cdot\begin{pmatrix}-1&0\\ 10&-1\end{pmatrix}.$
Both $A(0)$ and $A(1)$ correspond to global asymptotically stable systems, yet, for $s=\tfrac{1}{2}$, the system $\dot{x}=A(s)x$ is unstable. Hence, we cannot just construct straight-line homotopies between stable vector fields and expect that stability is preserved. Instead, we know from [ref:JongeneelSchwan2024TAC] that for vector fields with convex Lyapunov functions we should homotope via the canonical ODE $\dot{x}=-x$. Explicitly, consider the following path of vector fields defined by
$H(x;s):=\begin{pmatrix}-1&\max\{0,(1-2s)10\}\\ \max\{0,(2s-1)10\}&-1\end{pmatrix}x,\quad s\in[0,1].\penalty 9999\hbox{}\nobreak\hfill\quad\hbox{$\circ$}$
@@
[…]
@@ 1.2   Dynamical systems
We study semi-dynamical systems comprised of the triple $\Sigma:=({M}^{n},\varphi,\mathbb{R}_{\geq 0})$. Here, ${M}^{n}$ will be a smooth $n$-dimensional manifold (frequently, diffeomorphic to $\mathbb{R}^{n}$, which we denote by ${M}^{n}\simeq_{d}\mathbb{R}^{n}$,) and $\varphi:\mathbb{R}_{\geq 0}\times{M}^{n}\to{M}^{n}$ is a (global) semiflow, that is, a continuous map that satisfies for any $p\in{M}^{n}$: $(i)$ $\varphi(0,p)=p$; and $(ii)$ $(\varphi(s,\varphi(t,p))=\varphi(t+s,p)$ $\forall s,t\in\mathbb{R}_{\geq 0}:=\{t\in\mathbb{R}:t\geq 0\}$. We will usually write $\varphi^{t}(\cdot)$ instead of $\varphi(t,\cdot)$. In particular, we study (forward complete) semiflows generated by vector fields over ${M}^{n}$, that is, when $\varphi$ satisfies
@@
where $X$ is a continuous section, that is, the map $X:M^{n}\to TM^{n}$ is continuous and satisfies $\pi\circ X=\mathrm{id}_{M^{n}}$ for $\pi$ the projection $(p,v)\mapsto\pi(p,v)=p$. When $M^{n}=\mathbb{R}^{n}$, the identification $T\mathbb{R}^{n}\simeq_{d}\mathbb{R}^{n}\times\mathbb{R}^{n}$ allows for discussing vector fields as self-maps of $\mathbb{R}^{n}$. On $\mathbb{R}^{n}$, we will frequently work with the “canonical” ODE $\dot{x}=-x$, but also with the identity map $x\mapsto x$. To avoid confusion, we denote the canonical vector field by $-\partial_{x}$, and not $-\mathrm{id}_{\mathbb{R}^{n}}$.
The focus on semiflows instead of flows allows us to look at sufficiently regular discontinuous vector fields. This is relevant, as the introduction of feedback usually results in a closed-loop vector field that cannot be assumed to be continuous (e.g., think of optimal control [^2] One might also think of topological obstructions, however, the type of discontinuities we consider do not allow for overcoming those obstructions, in general [ref:ryan1994brockett].). This choice of setting is also motivated by other recent work. For instance, the research question in [ref:ozaslan2024exponential] is: “Given an exponentially stable optimization algorithm, can it be modified to obtain a finite/fixed-time stable algorithm?”. Some sufficient conditions are provided in [ref:ozaslan2024exponential], we will show a generic, yet less constructive viewpoint.
@@
where the set-valued map $F:\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$, with $2^{S}$ denoting the power set of a set $S$, is in some precise sense related to $X$. The rationale is to pass from a irregular single-valued map, to a more regular set-valued map.
@@ 1.3   Stability
In this section we will characterize stability under (1.2). Starting with the regular case, for simplicity, let the vector field $F$ be single-valued and smooth. In that case, $F$ generates a flow, denoted $\varphi(\cdot;F)$. A point $x^{\star}\in\mathbb{R}^{n}$ is an equilibrium point of $F$ when $F(x^{\star})=0$ or equivalently $\varphi^{t}(0;F)=0$ $\forall t\in\mathbb{R}$. We will set $x^{\star}:=0$, unless stated otherwise. Then, 00 is said to be globally asymptotically stable (GAS) (under $F$) when
@@
Although we merely work with $\mathbb{R}^{n}$ (or $M^{n}\simeq_{d}\mathbb{R}^{n}$) we take a topological approach akin to [ref:bhatiahajek2006local] and hence GAS is defined as above. One can turn Property (s.i) truly global, however, e.g., see [ref:wilson1969smoothing, ref:lin1996smoothV2]. Specifically, let $B_{r}(x)$ be some $r$-metric ball centered at $x$, then, require the existence of $\delta\in\mathcal{K}_{\infty}$ such that for any $\varepsilon\geq 0$ we have $\varphi^{t}(B_{\delta(\varepsilon)}(x))\subseteq B_{\varepsilon}(x)$ $\forall t\in\mathbb{R}_{\geq 0}$. In general these definitions are not equivalent, however, for 00 being GAS on $\mathbb{R}^{n}$, they are [ref:andriano1997global]. $\circ$
@@
Property (V.ii) implies sublevel set compactness. We call such a function a (smooth, strict and proper) Lyapunov function (with respect to the pair $(F,0)$). This work is about GAS, so we will omit “strict” and “proper” from now on. Then, based on converse theory, e.g., see [ref:Massera1956, ref:kurzweil1963inversion, ref:wilson1969smoothing, ref:fathi2019smoothing], we can appeal to the celebrated theorem stating that 00 is GAS if and only if there is a (corresponding) smooth Lyapunov function [ref:bacciotti2005liapunov, Thm. 2.4].
@@
One can check that $\varphi_{1}$ is a semiflow, describing a solution (e.g., in the sense of Filippov) to (1.3). Regarding stability, consider the $C^{\infty}$ Lyapunov function $x\mapsto V_{1}(x):=\tfrac{1}{2}x^{2}$ and see that $\nabla V_{1}(x)X_{1}(x)=-|x|<0$ on $\mathbb{R}\setminus\{0\}$ (consider a Huber loss to find a valid $W_{1}$ in the sense of Definition 1.3). This already shows that the existence of a smooth Lyapunov function, asserting that the origin is GAS, does not imply the existence of a flow, nor does it imply that convergence to 00 is merely asymptotic. Now let $X_{2}(x):=-\partial_{x}$, define $[0,1]\ni s\mapsto X(\cdot;s):=(1-s)X_{1}+sX_{2}$ and see that for and $s\in[0,1]$ we have $\nabla V(x)X(x;s)<0$ on $\mathbb{R}\setminus\{0\}$. The semiflow corresponding to $X(\cdot;s)$ becomes
@@
Then, $s\mapsto\varphi(\cdot;s)$ parametrizes a homotopy, along semiflows such that 00 remains GAS. For $\lim_{s\to 0^{+}}\varphi(\cdot;s)$, use $e^{-st}=1-st+\sum^{\infty}_{n=2}((-st)^{n}/n!)$. $\circ$
@@
It seems we cannot appeal to Filippov’s framework as $X_{3}$ is not locally essentially bounded. However, as $\nabla V_{\gamma}(x)=\gamma^{\prime}(\|x\|_{2})x/\|x\|_{2}$ we can study (1.5) directly. Decompose $x\in\mathbb{R}^{n}\setminus\{0\}$ as $x=\|x\|_{2}\cdot x/\|x\|_{2}=:r\cdot u$, then, $\dot{r}=-\gamma^{\prime}(r)$ while $\dot{u}=0$. In general $\gamma^{\prime}(s)>0$ $\forall s>0$, need not be true, but suppose this is true for our choice of $\gamma$, e.g., pick $s\mapsto\gamma(s)=s^{1/2}$. Now, suppose that $1/\gamma^{\prime}(s)$ is of class $\mathcal{K}$ on $\mathbb{R}_{\geq 0}$, we can define $\Gamma(r):=\int^{r}_{0}1/\gamma^{\prime}(\rho)\mathrm{d}\rho$, which is now of class $\mathcal{K}_{\infty}$ and hence invertible on $\mathbb{R}_{\geq 0}$. Under the aforementioned assumptions, the semiflow corresponding to $X_{3}$ becomes
$(t,x)\mapsto\varphi_{3}^{t}(x):=\begin{cases}\min\{0,\Gamma^{-1}(\Gamma(\|x\|_{2})-t)\}x/\|x\|_{2}\quad&\Gamma^{-1}(t)\leq\|x\|_{2}\\ 0\quad&\mathrm{otherwise}\end{cases}.\penalty 9999\hbox{}\nobreak\hfill\quad\hbox{$\circ$}$

### v3

@@ 1  Introduction
In the context of what we call today Conley index theory (see Section A), Conley posed the following “converse question” in the 1970s: “To what extent does the homotopy index [Conley index] itself determine the equivalence class of isolated invariant sets which are related by continuation?[ref:conley1978isolated, p. 83]. Recently, Kvalheim proved that uniquely integrable $C^{0}$ vector fields, on a $C^{\infty}$ manifold $M$, rendering a compact set $A\subseteq M$ asymptotically stable, are homotopic on an open neighbourhood $U\supseteq A$ such that throughout the homotopy the vector fields do not vanish on $U\setminus A$ [ref:kvalheim2022obstructions, Thm. 1]. Connecting this result to Conley, a follow-up question (revitalized) by Kvalheim—which is the central question of this note—is the following.
@@
This question, in one form or another, inspired several works, for instance, [ref:reineck1992continuation, ref:mrozek2000conley, ref:JongeneelSchwan2024TAC]. Here, we will elaborate on commentary by the author in [ref:JongeneelSchwan2024TAC]. Specifically, in the seminal paper “Asymptotic stability equals exponential stability, and ISS equals finite energy gain—if you twist your eyes” from 1999, Grüne, Sontag and Wirth showed that asymptotic stability “equals” exponential stability in the sense that if an equilibrium point is asymptotically stable under some vector field on $\mathbb{R}^{n}$, then, there is a suitable change of coordinates rendering this point exponentially stable [ref:grune1999asymptotic]. Such a change of coordinates is understood to be instantaneous. However, by leveraging their work, we show that asymptotic stability can be continuously “transformed” into exponential stability, while preserving asymptotic stability throughout the transformation, see Theorem 3.6. Differently put, asymptotic stability equals exponential stability—not only if you twist your eyes, but while you twist your eyes. The key to this is to observe that we can select the transformation from [ref:grune1999asymptotic] to be an orientation-preserving homeomorphism on $\mathbb{R}^{n}$, not just any homeomorphism cf. [ref:JongeneelSchwan2024TAC, Sec. III], see the proof of Proposition 3.2.
@@ 1.1  Related work
It can be argued that questions of the form as above, emerged from studies aimed at classifying manifolds, maps and so forth. A successful, yet coarse, resolution has been found in the study of these objects up to homotopy, e.g., motivated early on by the fundamental group being homotopy invariant, Hopf’s degree theorem, CW complexes, intractability of topological equivalence and more at the intersection of topology and dynamical systems.
@@
Besides possibly stronger necessary conditions for continuous stabilization, the study of Conley’s converse question directly relates to understanding topological properties of spaces of stable dynamical systems (as was motivating prior work [ref:JongeneelSchwan2024TAC]). This is of practical importance as several frameworks in optimal control and reinforcement learning aim to optimize over precisely such a space, e.g., moving from a pre-stabilized system to an optimally controlled system, we point to [ref:Fazel_18, ref:furieri2024learning] and references therein.
@@
$0,1]\ni s\mapsto A(s):=s\cdot\begin{pmatrix}-1&10\\ 0&-1\end{pmatrix}+(1-s)\cdot\begin{pmatrix}-1&0\\ 10&-1\end{pmatrix}$.
Both $A(0)$ and $A(1)$ correspond to global asymptotically stable systems, yet, for $s=\tfrac{1}{2}$, the system $\dot{x}=A(s)x$ is unstable. Hence, we cannot simply construct straight-line homotopies between stable vector fields and expect that stability is preserved. Instead, we know from [ref:JongeneelSchwan2024TAC] that for vector fields with convex Lyapunov functions we should homotope via the canonical ODE $\dot{x}=-x$. Explicitly, consider the following path of vector fields defined by
$H(x;s):=\begin{pmatrix}-1&\max\{0,(1-2s)10\}\\ \max\{0,(2s-1)10\}&-1\end{pmatrix}x,\quad s\in[0,1].\penalty 9999\hbox{}\penalty 10000\hfill\quad\hbox{$\circ$}$
@@
For $r\in\mathbb{N}\cup\{\infty\}$, $C^{r}(U;V)$ denotes the set of $C^{r}$-smooth functions from $U$ to $V$. The inner product on $\mathbb{R}^{n}$ is denoted by $\langle\cdot,\cdot\rangle$ and $\mathbb{S}^{n-1}:=\{x\in\mathbb{R}^{n}:\|x\|_{2}=1\}$. The identity map $p\mapsto p$ on a space $M$ is denoted by $\mathrm{id}_{M}$. Given a function $V:M\to\mathbb{R}$, then, $V^{-1}(c):=\{p\in M:V(p)=c\}$. We typically use $\varphi$ to denote a (semi)flow. If this flow comes from a vector field $X$, we write $\varphi(\cdot;X)$, similarly, if $\varphi$ is parametrized by $s\in[0,1]$ we write $\varphi(\cdot;s)$, that is, we overload the meaning of $\varphi$. A homotopy $0,1]\times M\ni(s,p)\mapsto H(s,p)$ is said to be an isotopy when $p\mapsto H(s,p)$ is a topological embedding for all $s\in[0,1]$. A homeomorphism $\psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be a stable homeomorphism when it is a finite composition of homeomorphisms that equal the identity map on some—not necessarily the same—non-empty open subset of $\mathbb{R}^{n}$. We denote by $\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$ the group of homeomorphisms $\psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$. Additionally, we denote by $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})\subset\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$ the subgroup of all orientation-preserving [^1] Regarding orientation, we point to [ref:guillemin2010differential, Lee2] for the smooth case. The topological definition relies on algebraic topology, see [Hatcher, Sec. 3.3]. homeomorphisms. A continuous function $\alpha:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is of class $\mathcal{K}$ when $\alpha$ is strictly increasing and $\alpha(0)=0$. If, additionally $\lim_{s\to+\infty}\alpha(s)=+\infty$, then, $\alpha$ is of class $\mathcal{K}_{\infty}$. At last, a continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is said to be of class $\mathcal{KL}$ when $s\mapsto\beta(s,t)$ is of class $\mathcal{K}$ and $t\mapsto\beta(s,t)$ is decreasing and such that $\lim_{t\to+\infty}\beta(s,t)=0$.
@@ 1.2  Dynamical systems
We study semi-dynamical systems comprised of the triple $\Sigma:=({M}^{n},\mathbb{R}_{\geq 0},\varphi)$. Here, ${M}^{n}$ will be a smooth $n$-dimensional manifold (frequently, diffeomorphic to $\mathbb{R}^{n}$, which we denote by ${M}^{n}\simeq_{d}\mathbb{R}^{n}$,) and $\varphi:\mathbb{R}_{\geq 0}\times{M}^{n}\to{M}^{n}$ is a (global) semiflow, that is, a continuous map that satisfies for any $p\in{M}^{n}$: $(i)$ $\varphi(0,p)=p$; and $(ii)$ $(\varphi(s,\varphi(t,p))=\varphi(t+s,p)$ $\forall s,t\in\mathbb{R}_{\geq 0}:=\{t\in\mathbb{R}:t\geq 0\}$. We will usually write $\varphi^{t}(\cdot)$ instead of $\varphi(t,\cdot)$. In particular, we study (forward complete) semiflows generated by vector fields over ${M}^{n}$, that is, when $\varphi$ satisfies
@@
where $X$ is a continuous section, that is, the map $X:M^{n}\to TM^{n}$ is continuous and satisfies $\pi\circ X=\mathrm{id}_{M^{n}}$ for $\pi$ the projection $(p,v)\mapsto\pi(p,v)=p$. When $M^{n}=\mathbb{R}^{n}$, the identification $T\mathbb{R}^{n}\simeq_{d}\mathbb{R}^{n}\times\mathbb{R}^{n}$ allows for discussing vector fields as self-maps of $\mathbb{R}^{n}$. On $\mathbb{R}^{n}$, we will frequently work with the “canonical” ODE $\dot{x}=-x$, but also with the identity map $x\mapsto x$. To avoid confusion, we denote the canonical vector field by $\partial_{x}$, and not $\mathrm{id}_{\mathbb{R}^{n}}$.
The focus on semiflows instead of flows allows us to look at sufficiently regular discontinuous vector fields. This is relevant, as the introduction of feedback usually results in a closed-loop vector field that cannot be assumed to be continuous (e.g., think of optimal control [^2] One might also think of topological obstructions, however, the type of discontinuities we consider do not allow for overcoming those obstructions, in general [ref:ryan1994brockett].). This choice of setting is also motivated by other recent work. For instance, the research question in [ref:ozaslan2024exponential] is: “Given an exponentially stable optimization algorithm, can it be modified to obtain a finite/fixed-time stable algorithm?”. Some sufficient conditions are provided in [ref:ozaslan2024exponential], we will show a generic, yet less constructive, viewpoint.
@@
where the set-valued map $F:\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$, with $2^{S}$ denoting the power set of a set $S$, is in some precise sense related to $X$. The rationale is to pass from an irregular single-valued map, to a more regular set-valued map.
@@ 1.3  Stability
In this section we will characterize stability under (1.2). Starting with the regular case, for simplicity, let the vector field $F$ be single-valued and smooth. In that case, $F$ generates a flow, denoted $\varphi(\cdot;F)$. A point $x^{\star}\in\mathbb{R}^{n}$ is an equilibrium point of $F$ when $F(x^{\star})=0$ or equivalently $\varphi^{t}(x^{\star};F)=x^{\star}$ $\forall t\in\mathbb{R}$. We will set $x^{\star}:=0$, unless stated otherwise. Then, 0 is said to be globally asymptotically stable (GAS) (under $F$) when
@@
Although we merely work with equilibria and $\mathbb{R}^{n}$ (or $M^{n}\simeq_{d}\mathbb{R}^{n}$) we take a topological approach akin to [ref:bhatiahajek2006local] and hence GAS is defined as above. One can turn Property (s.i) truly global, however, e.g., see [ref:wilson1969smoothing, ref:lin1996smoothV2]. Specifically, let $B_{r}(x)$ be some $r$-metric ball centered at $x$, then, require the existence of $\delta\in\mathcal{K}_{\infty}$ such that for any $\varepsilon\geq 0$ we have $\varphi^{t}(B_{\delta(\varepsilon)}(x))\subseteq B_{\varepsilon}(x)$ $\forall t\in\mathbb{R}_{\geq 0}$. In general these definitions are not equivalent, however, for 0 being GAS on $\mathbb{R}^{n}$, they are [ref:andriano1997global].  $\circ$
@@
Property (V.ii) implies sublevel set compactness. We call such a function a (smooth, strict and proper) Lyapunov function (with respect to the pair $(F,0)$). This work is about GAS, so we will omit “strict” and “proper” from now on. Then, based on converse theory, e.g., see [ref:Massera1956, ref:kurzweil1963inversion, ref:wilson1969smoothing, ref:fathi2019smoothing], we can appeal to the celebrated theorem stating that 0 is GAS if and only if there is a (corresponding) smooth Lyapunov function [ref:bacciotti2005liapunov, Thm. 2.4].
@@
One can check that $\varphi_{1}$ is a semiflow, describing a solution (e.g., in the sense of Filippov) to (1.3). Regarding stability, consider the $C^{\infty}$ Lyapunov function $x\mapsto V_{1}(x):=\tfrac{1}{2}x^{2}$ and see that $\nabla V_{1}(x)X_{1}(x)=-|x|<0$ on $\mathbb{R}\setminus\{0\}$. This already shows that the existence of a smooth Lyapunov function, asserting that the origin is GAS, does not imply the existence of a flow, nor does it imply that convergence to 0 is merely asymptotic. Now let $X_{2}(x):=-\partial_{x}$, define $0,1]\ni s\mapsto X(\cdot;s):=(1-s)X_{1}+sX_{2}$ and see that for and $s\in[0,1]$ we have $\nabla V(x)X(x;s)<0$ on $\mathbb{R}\setminus\{0\}$. The semiflow corresponding to $X(\cdot;s)$ becomes
@@
Then, $s\mapsto\varphi(\cdot;s)$ parametrizes a homotopy, along semiflows such that 0 remains GAS. For $\lim_{s\to 0^{+}}\varphi(\cdot;s)$, use $e^{-st}=1-st+\sum^{\infty}_{n=2}((-st)^{n}/n!)$.  $\circ$
@@
It seems we cannot appeal to Filippov’s framework as $X_{3}$ is not locally essentially bounded. However, as $\nabla V_{\gamma}(x)=\gamma^{\prime}(\|x\|_{2})(x/\|x\|_{2})$ we can study (1.5) directly. Decompose $x\in\mathbb{R}^{n}\setminus\{0\}$ as $x=\|x\|_{2}\cdot(x/\|x\|_{2})=:r\cdot u$, then, $\dot{r}=-\gamma^{\prime}(r)$ while $\dot{u}=0$. In general $\gamma^{\prime}(s)>0$ $\forall s>0$, need not be true, but suppose this is true for our choice of $\gamma$, e.g., pick $s\mapsto\gamma(s)=s^{1/2}$. Now, suppose that $s\mapsto 1/\gamma^{\prime}(s)$ is of class $\mathcal{K}$ on $\mathbb{R}_{\geq 0}$, we can define $r\mapsto\Gamma(r):=\int^{r}_{0}1/\gamma^{\prime}(\varrho)\mathrm{d}\varrho$, which is now of class $\mathcal{K}_{\infty}$ and hence invertible on $\mathbb{R}_{\geq 0}$. Under the aforementioned assumptions, the semiflow corresponding to $X_{3}$ becomes
$(t,x)\mapsto\varphi_{3}^{t}(x):=\begin{cases}\min\{0,\Gamma^{-1}(\Gamma(\|x\|_{2})-t)\}(x/\|x\|_{2})\quad&\Gamma^{-1}(t)\leq\|x\|_{2}\\ 0\quad&\mathrm{otherwise}\end{cases}.\penalty 9999\hbox{}\penalty 10000\hfill\quad\hbox{$\circ$}$